ou aro t ng two courses this sernester tology and chemistry Vou have quines comg upinboth classe. The table below sho s your grade on ead ar kran ont t les oton-h implies that one hour of studying for Chemistry will generate an expected grade of 2 on Chemistry, whereas one hour of studying for Biology will ganerale an expected grade of 82 on Blogy ﹃fura h Far-m.oM Use the dea of opinzaton in dt wees ๒ decide how much 5me you would spend storingfor earnquiz ,you had a bb/d 1,2.or3tonnp kve If you had 2 hours, then you should study houra) for chemistry and hourt tor bioogy Ento our s
Answer #1

If the person has only one hour to study, he should choose to study Chemistry as the total marks by studing 1 hour chemistry and 0 hours biology = 62+ 56=118 which is greater than the score that will be obtained by studying 1 hour biology and 0 hours chemistry.

If the person has two hours , he has three senarios

a) 2 hours chemistry and 0 hours biology: Total score= 67+56 = 123

b) 2 hours biology and 0 hours chemistry: Total score= 55+66 = 121

c) 1 hour biology and 1 hour chemistry: Total score = 62+62=124

So the person should study one 1 hour for chemistry and 1 hour for biology.

If the person had 3 hours , he has two options:

a) 2 hours chemistry and 1 hour biology: Total score= 67+62= 129

b) 2 hours biology and 1 hour chemistry: Total score= 62+66 = 128

So if the person has 3 hours , he should study 2 hours for chemistry and 1 hour for biology.